OpenStudy (anonymous):

Jimmy has seven times as many dimes as nickels. Their value is $3.75. How many nickels does he have? A.35 B.15 C.5

OpenStudy (anonymous):

no il give you best anwser for the anwser

OpenStudy (anonymous):

10(7n) + 5n = 375 70n + 5n = 375 75n = 375 n = 5 (# of nickels) d = 7n*7*5 = 35